OpenStudy (anonymous):

The following table shows the probability of rolling the numbers 1 through 7 on a fair seven-sided die. X 1 2 3 4 5 6 7 P(X) 1/7 1/7 1/7 1/7 1/7 1/7 1/7 What is the standard deviation of the random variable X, "the number that comes up on the die?"

OpenStudy (anonymous):

Do you know the standard deviation formula?

OpenStudy (anonymous):

Is this the question you pinged about?

OpenStudy (anonymous):

If you look here: http://en.wikipedia.org/wiki/Binomial_distribution you will see that the variance = np(1-p) Does this help?

OpenStudy (anonymous):

ok first of all you will need mean of the above given by \[\Large \mu= \sum_{i=1}^{n}X_{i}.P_{i}(x)\] this means multiply the random variable with its corresponding probability and add them. can yu find mean??

OpenStudy (anonymous):

@lilimsan find mean and let me know !

OpenStudy (anonymous):

@sami-21 is this not a binomial distribution question . Can't we assume n p (1-p) for variance?

OpenStudy (anonymous):

@telliott99 in binomial you have either yes or no. in this case it is easy to use discrete random variable. yes binomial is also discrete distribution and if you know both probability of both success and failure , use it.

OpenStudy (anonymous):

Have to think about it. So the issue is X, "the number that comes up on the die?" @sami-21

OpenStudy (anonymous):

@lilimsan I think that for binomial distribution with p = 1/7 say a 7-sided die labeled 1 through 7 where the P of a particular result is 1/7 that the variance for that in a number of trials (say n = 100) will be n p (1-p) = 100 1/7 (6/7) And the standard deviation will be the square root of that.